40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8385     Entry time: Mon Apr 1 17:59:04 2013
Author: Jenne 
Type: Update 
Category: Locking 
Subject: PRMI videos 

[Gabriele's work, I'm just spectating]

Annalisa is working on finding the PSL/AUX laser beatnote, so the PSL temp is changing, but Gabriele is still able to lock.  Here are some videos:



ELOG V3.1.3-